OpenStudy (anonymous):

You drop a rock off a bridge. The rock’s height, h (in feet above the water), after t seconds is modeled by h = -16t2 + 541. What is the height of the rock after 2 seconds? A. –64 feet B. 605 feet C. 509 feet D. 477 feet

OpenStudy (anonymous):

It's easy! Substitute 2 for (t) so h= -16(2)^2+541. Do parenthesis first and then exponents 2^2 = 4. h=-16(4)+541. Now do the multiplication -16*4= -64 h=-64+541...h = 477

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Well, I hope this helped!

OpenStudy (anonymous):

Sorry reading. Im dexlic

OpenStudy (anonymous):

Oh. That's okay :)

OpenStudy (anonymous):

Well, the answer is 477!

OpenStudy (anonymous):

Thank you ^3^

OpenStudy (anonymous):

:3

OpenStudy (anonymous):

Anytime!